OpenStudy (thaliafortunx):

What is the sum of the geometric series in which a1 = 7, r = 3, and an = 1,701? Hint: Sn= a1(1-rn)/1-r, r ≠ 1, where a1 is the first term and r is the common ratio.

OpenStudy (callie_love_me):

That formula basically gives it out. Just plug the numbers into the Sn formula.

OpenStudy (mhchen):

So you have the geometric series formula: \[S_{n} = a_{1} * \frac{ 1-r^{n} }{ 1-r }\] a1 = 7 r = 4 I hope you know an = a1 * r^n so an = 7 * (4^n) an/7 = 4^n \[\log _{4}(1701) = n\] And plug a1, r, and n into that formula at the top.